OpenStudy (anonymous):

what is cot2a=0 Thus 2a=90˚....how did this happen?

OpenStudy (mimi_x3):

\[ cot2a = 0 => \frac{cos2a}{sin2a} = 0 => cos2a = 0 => 2a = cos(0)\]

OpenStudy (mimi_x3):

wait i made a typo: \[ 2a = cos^{-1}(0)\]

OpenStudy (mimi_x3):

since you know that \(cos^{-1}(0) = 90\)

OpenStudy (mimi_x3):

then \( 2a = 90\)

OpenStudy (anonymous):

wait how did cot2a= cos2a/sin2a?

OpenStudy (anonymous):

Firstly you should remember that: \[\cos(90) = 0\]

OpenStudy (mimi_x3):

because \[ cota = \frac{cosa}{sina}\]

OpenStudy (anonymous):

*cot

OpenStudy (anonymous):

\[\large \cot(90^{\circ}) = 0\]

OpenStudy (anonymous):

So you can replace 0 with cot(90): \[\cot(2a) = \cot(90)\]

OpenStudy (anonymous):

im sorry but isnt cot= 1/tana?

OpenStudy (mimi_x3):

well you can use that method if you remember whats cot(90)

OpenStudy (anonymous):

Now as trigonometric ratio is same: So you can equate the angles: \[2a = 90\]

OpenStudy (mimi_x3):

yes it because what tana?

OpenStudy (mimi_x3):

what is*

OpenStudy (mimi_x3):

yes it is but whats tan(a) then?*

OpenStudy (anonymous):

oh cot also equals to x/y so cos/sin right??

OpenStudy (anonymous):

Right..

OpenStudy (anonymous):

ohhhh okay it makes sense

OpenStudy (anonymous):

so i get that cos2a/sin2a = 0 but then what is next?

OpenStudy (mimi_x3):

i typed the solution for you already..

OpenStudy (anonymous):

Multiply by sin(2a) both the sides..

OpenStudy (anonymous):

so it becomes cos2a (sin2a) = 0 (sin2a)

OpenStudy (anonymous):

No.. There is one sin(2a) in the denominator too..

OpenStudy (anonymous):

Can't you cancel sin(2a) with sin(2a) on left hand side ??

OpenStudy (anonymous):

so does it become like this?

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

oh okay now i get the steps! thank you so much @waterineyes @Mimi_x3

OpenStudy (anonymous):

Welcome dear..